OpenStudy (anonymous):

Agh can someone help pls 2+log5 of 10 + log5 of 2 - 5 Having a blank moment :(

OpenStudy (anonymous):

Well, for starters we can simplify 2-5 (I'm sure you can do that easily) and then we look at \[\log_{5} 10 + \log_{5}2 \] Using a certain log law, you could turn that into one log expression.

OpenStudy (anonymous):

so i get -3log5 of 20??

OpenStudy (anonymous):

Is your question in the form \[2 + \log_{5}10 + \log_{5}2 - 5? \] If so, the -3 and the log 5 of 20 would be two separate terms: \[-3 + \log_{5}20 \] or \[\log_{5}20 - 3\]

OpenStudy (anonymous):

ah ok and no further simplification?

OpenStudy (anonymous):

I don't think so - you could try it though. Using the change of base rule, you could put \[\frac{ \ln 20 }{ \ln5 }\] to evaluate the log, then subtract 3, but your answer will just be a decimal approximation. If I was you, I would try it in all cases anyway, because sometimes it evaluates to an exact integer, which could be useful.

OpenStudy (anonymous):

For example, if you had \[\log_{4}16 \] and didn't immediately see what that could evaluate to, you could check it out using the change of base rule and discover it evaluates to 2.

OpenStudy (anonymous):

\[\log_{3}x^5+\log_{3}x^2-\log_{3}x^4 \]

OpenStudy (anonymous):

is the answer log3x^6??

OpenStudy (anonymous):

\[\log_{3}x^6 \]

OpenStudy (anonymous):

Consider what you're doing here. When you add the logs to turn them into one log, you're multiplying the inside of the logs together (and when you're subtracting, you divide by the inside of the log). In this case, you have \[\log_{3} \frac{ x^5 \times x^2 }{ x^4 }\]

OpenStudy (anonymous):

What should that evaluate to be?

OpenStudy (anonymous):

\[\log_{3}x^3 \]

OpenStudy (anonymous):

That looks correct to me! You could also use the rule that \[\log_{b} x^n = nlog_{b}x\] to get 3 outside of log 3 of x, as your answer.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

\[3^{2x}-12*3^{x}+27=0\]

OpenStudy (anonymous):

This type of question involves a substitution. I'll give you the general structure to work these out. 1. Let u = *something that can be represented as a variable of a quadratic* 2. Solve the quadratic in terms of u. 3. Allow the u to become that *something* again, and you'll usually have a couple of equations. 4. Solve these equations for x, if possible.

OpenStudy (anonymous):

ok thanks for that what about 3^x = 30 solve for x

OpenStudy (anonymous):

Use the change of base rule: \[b^x = y\] \[x = \frac{ \ln y }{ \ln b }\]

OpenStudy (anonymous):

A way I like to remember it is "base on the bottom", in case you forget which number goes on the bottom and which goes on the top.

OpenStudy (anonymous):

I am still lost on the previous one Would I say let 3^2x = y??? and therefore 3^x also = y

OpenStudy (anonymous):

or is it 3^x = y??

OpenStudy (anonymous):

By previous one, do you mean the one involving a substitution and quadratic? If so, consider \[ 3^{2x} = \left( 3^x \right)^2\]

OpenStudy (anonymous):

so therefore if I say let 3x = y then I would get y^3 - 12 =27?

OpenStudy (anonymous):

If you let 3x = y, you would get y^2 -12y + 27 = 0, and you could then solve that as a quadratic.

OpenStudy (anonymous):

ahhhh ok thanks so much :)

OpenStudy (anonymous):

\[\log_{4} (0.5x+1)=2\] how do I go about this?

OpenStudy (triciaal):

4^2 = 0.5x + 1 solve for x

OpenStudy (anonymous):

so new to all of this :( Do you ever get the hang of it?

OpenStudy (triciaal):

it gets very easy with practice

OpenStudy (anonymous):

\[\log_{2}x+\log_{2}5=2 \]

OpenStudy (triciaal):

indices, logs and surds one of the fun sections of math I personally don't care for statistics i have to really pay attention then

OpenStudy (anonymous):

lol

OpenStudy (triciaal):

add logs multiply the numbers

OpenStudy (triciaal):

so 2^2 = 5 x x =

OpenStudy (anonymous):

\[\log_{2}5x=2 \] how do i isolate the x?

OpenStudy (triciaal):

see above

OpenStudy (anonymous):

x=4/5

OpenStudy (triciaal):

you got it already see how easy

OpenStudy (triciaal):

just need to know the log laws

OpenStudy (anonymous):

they are so simple but so easy to forget :(

OpenStudy (triciaal):

be careful when converting bases

OpenStudy (anonymous):

\[\log_{6}(x-2)+\log_{6}(x+3)=1 \]

OpenStudy (triciaal):

hint log base 6 of 6 = 1

OpenStudy (anonymous):

also I put this up before but I have to solve for x and can't remember where I wrote it \[3^{x}=30\]

OpenStudy (triciaal):

what do you have for the first one?

OpenStudy (anonymous):

(x-2) (x+3)=6?

OpenStudy (triciaal):

yes keep going

OpenStudy (triciaal):

x log base 10 of 3 = log base 10 of 30 x = log base 10 of 30/ log base 10 of 3

OpenStudy (triciaal):

are you allowed to use the calculator?

OpenStudy (anonymous):

yes

OpenStudy (triciaal):

well you can do the 2nd one like that did you get 4, -3 for the other one

OpenStudy (anonymous):

\[x^{2}-5x=0???\]

OpenStudy (anonymous):

oooo I am a bit lost LOL

OpenStudy (triciaal):

(x-2) (x+3)=6 x(x + 3) -2(x + 3) = 6 x^2 + 3x -2x -6 = 6

OpenStudy (anonymous):

agh I didn't have a plus on my 3.... so (x-4) (x+3) = 0

OpenStudy (triciaal):

so 4, -3

OpenStudy (anonymous):

would it be true to say that -3 is not a solution as you can't have a log of a negative?

OpenStudy (triciaal):

i didn't check in the original. always a good practice

OpenStudy (triciaal):

same page

OpenStudy (anonymous):

it looks like 4 is not a solution??

OpenStudy (triciaal):

why

OpenStudy (anonymous):

i think I am calculating wrong ahhhh I am a numbskull :(

OpenStudy (triciaal):

log14/log 6 =1.14/.77=1.48 log6(x−2)+log6(x+3)=1 log base 6 of (4-2)*(4 + 3) 2*7 =log base 6 of 14

OpenStudy (anonymous):

so it doesn't equal 1?

OpenStudy (triciaal):

log base 6 of 14 = 1.48 which is almost 1

OpenStudy (anonymous):

ok i think lol

OpenStudy (anonymous):

oh dear I just had a look it is (x-2) + (x+3) not multiply :( @triciaal

OpenStudy (triciaal):

log6(x−2)+log6(x+3)=1 when we add logs we multiply the numbers in this case expressions ? why do you say not multiply?

OpenStudy (anonymous):

oh so it still becomes (x-2) (x+3) = 1 you must think I am a twat LOL